# NeuSemSlice: Towards Effective DNN Model Maintenance via Neuron-level Semantic Slicing

### v1

@@ 5. Evaluation
In this section, we evaluate the performance of NeuSemSlice. We first outline the experimental setup and then, we introduce our evaluation aiming to answer the following research questions:
@@
  • RQ2: Is NeuSemSlice effective for model restructuring?
  • RQ3: Is NeuSemSlice effective for model re-adaption?
@@ 5.1. Evaluation Setup
To ensure the comprehensiveness and depth of the experiments, we select four datasets widely utilized in the deep neural network domain, along with five network models for experimental evaluation. ❶ MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017). For training, we use a Multilayer Perceptron (MLP) model with two hidden layers, containing 512 and 256 neurons respectively. ❷ CIFAR-10 (Krizhevsky et al., 2009). Three deep neural network models—AlexNet (Krizhevsky et al., 2012), VGG-16 (Simonyan and Zisserman, 2014), and ResNet-101 (He et al., 2016)—are utilized for training. ❸ ImageNet (Deng et al., 2009). Following the setup in (Xie et al., 2022), only data from the top 10 categories are selected for experiments, with the VGG-19 model employed.
@@
All these experiments are conducted on the Intel Xeon Silver 4214 Processor with 2 Tesla V100 GPUs with 32GB memory. We have implemented our tool based on Pytorch (Paszke et al., 2019a). The source code, related information of datasets and models, and complete experimental results are released at (NeuSemSlice, 2024).
@@ 5.2. RQ1: The performance of semantic slicing
Detailed Configurations. In RQ1, we compare the following three slicing methods: global slicing, category-wise slicing, and semantic slicing. To assess neuron contribution and select critical neurons, we utilize five neuron contribution metrics: Avg, Var, Ens, DeepLIFT, and Taylor. As outlined in Section 3.1, global slicing and category-wise slicing choose top 10%, 20%, … up to 100% of neurons in each layer or category for slicing respectively. Additionally, we also introduce DeepArc for comparative analysis. Following the settings used in (Ren et al., 2023), we set the threshold at 0.99 to identify redundant layers in the model. Subsequently, we remove these redundant layers and compare the model accuracy when only critical semantic layers are retained.
@@
Comparison with Global Slicing and Category-wise Slicing. Figure 5 compares the slicing precision differences among semantic slicing, global slicing, and category-wise slicing (The complete experimental results can be found in (NeuSemSlice, 2024)). For complex models like AlexNet and VGG16, global slicing and category-wise slicing can only maintain certain model performance at very low compression rates, with accuracy significantly declining as the compression rate increases. Semantic slicing demonstrates significant improvement over the other two techniques across all five neuron contribution metrics. For example, in the AlexNet-CIFAR10 case (Figure 5(c)), semantic slicing uses the neuron contribution metric Taylor to maintain model performance, with only a 0.38% loss in accuracy at a 50% compression rate. Conversely, the other two techniques experience significant accuracy losses of around 70% at the same compression rate. Similarly, on VGG16 using the DeepLIFT metric (Figure 5(d)), semantic slicing at around a 10% compression rate reduces accuracy loss by 41.18% and 77.77% compared to global slicing and category-wise slicing.
@@
These results adequately confirm that inadequate account for the uneven distribution of critical neurons across different categories and network layers would lead to poor slicing performance. This highlights the advantage of the proposed technique, which can dynamically identify critical features and components of the model based on semantic similarity.
The Effects of DeepArc. For DeepArc, we identify redundant layers in six settings based on layer similarity, as shown in Figure 5. Firstly, for the MLP, the model only has two hidden layers. When any layer is removed, we follow the model restructuring detailed in DeepArc and introduce a “sewing layer” to align the input and output of the layers. However, this results in no change in the size of the model, indicating that DeepArc cannot effectively identify redundancies in such models. In the other four configurations, DeepArc only identified redundant layers with a similarity greater than 0.99 in VGG16 and ResNet101.
@@
5.3. RQ2: Model Restructuring
Table 4. Results of Model Restructuring.
@@ 5.3. RQ2: Model Restructuring
Detailed Configuration. To systematically assess NeuSemSlice for the model restructuring task, we adopt the same experimental setup detailed in (Ren et al., 2023), and validate by comparing the compression efficiency and effects on the model before and after applying NeuSemSlice. Specifically, after removing the non-critical neurons of the model, we expect the compression technology to directly focus on the redundant parts within the critical neurons, improving the iteration efficiency [^3] It is important to note that in this experiment and the subsequent RQ3 and RQ4, our method is based on the optimal configuration identified in RQ1. The term “our method combined with DeepArc” refers to retraining the critical neurons identified by our method in the critical modules recognized by DeepArc (targeting adversarial samples).. Based on this, we apply four of the most advanced model compression techniques currently available—LAMP (Lee et al., 2020), Global (Morcos et al., 2019), Uniform+ (Gale et al., 2019), and ERK (Evci et al., 2019)—for optimal compression results.
@@
Results and Findings. Table 4 presents the experimental results of NeuSemSlice in model restructuring. Across all experimental settings, the application of NeuSemSlice in model restructuring allows models to achieve higher compression ratios with fewer iterations, indicating that NeuSemSlice can improve both compression efficiency and effectiveness. Apart from the VGG19-ImageNet configuration, NeuSemSlice not only enhances compression efficiency and rates but also slightly improves model accuracy. Specifically, in six experimental settings, the average number of iterations required for model compression decreases from 8.67 to 7.50, enhancing efficiency by 13.46%, and the model compression ratio increases from 37.26% to 40.26%. These results demonstrate that the application of NeuSemSlice in model restructuring boosts the performance of model compression and enhances the performance of compressed models.
Answer to RQ2: For model restructuring, NeuSemSlice not only elevates the efficiency of compression technology—an average increase of 13.46%—but also achieves approximately a 3% enhancement in the compression rate and a slight increase in accuracy.
@@ 5.4. RQ3: Model Re-adaptation
Table 5. Results of Model Readaptation.
@@
Method TPR $\downarrow$ RA $\uparrow$ RAA TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$
@@
Detailed Configuration. To answer RQ3, we conduct an experiment using two prevalently-employed adversarial sample generation techniques, Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014) and Projected Gradient Descent (PGD) (Madry et al., 2017). We generate 1,000 adversarial samples, with FGSM reducing the model’s average accuracy to 18.53%, and PGD to 3.8%. We aim to compare the Training Parameter Ratio (TPR) and the Retrained Accuracy (RA) of existing approaches and our proposed method. Specifically, the experiment settings include fully training (i.e., comparison baseline), DeepArc, NeuSemSlice, and a combination of NeuSemSlice with DeepArc under identical training epochs. Additionally, this RQ further assesses the resilience of each method against similar adversarial attacks. We generate another 1,000 adversarial samples after the model re-adaption, and obtain the model accuracy against the samples (i.e., shown as Re-Attack Accuracy, RAA).
Results and Findings. Table 5 presents our experimental results in the four settings. Compared to the traditional method of fully training) and DeepArc, our method demonstrates a significant reduction in the required model parameters for training, averaging only 61.49%. This represents an average decrease of 38.51% in training parameters compared to the fully training strategy and an average reduction of 16.66% reduction relative to DeepArc. Moreover, our method surpasses fully training by an average of 0.67% and DeepArc by an average of 0.50% in Retrained Accuracy, demonstrating advantage in both training cost-efficiency and effectiveness. Among the four methods, our method combined with DeepArc achieves a 0.73% increase in Retrained Accuracy compared to traditional training methods, with the lowest training cost (averaging only 47.75% of the parameters trained).
Our method combined with DeepArc also shows the best performance in terms of re-attack performance, achieving up to a 17.8% improvement in RAA. The results further demonstrate the superior performance of our method in model re-adaptation scenarios. However, we observe that all of the methods experience notable performance loss against re-attack, alerting us to design better methods against adversarial attacks in the future, to be discussed in Section 5.6. Finally, NeuSemSlice is comparable to DeepArc in efficiency, requiring only 93.53% of the time needed for full training on average. When NeuSemSlice is combined with DeepArc, it requires only 89.80% of the full training time. The detailed data is available at (NeuSemSlice, 2024).
@@ 5.5. RQ4: Model Incremental Development
Table 6. Results of Model Incremental Development.
@@
Detailed Configuration. To answer RQ4, we split the datasets MNIST, FMNIST, CIFAR10, and ImageNet into two subtasks, each with five categories, labeled as task $A$ and task $B$. For instance, in the CIFAR10 dataset, which contains ten distinct image categories, task $A$ consists of the first five categories, while task $B$ consists of the remaining five categories.
@@
To simulate the process of incremental model development, we start by training models exclusively on task $A$ in six different experimental setups following the division. Then, we proceed to train these models on task $B$ while aiming to preserve their performance on task $A$. The evaluation focuses on the accuracy for both task $A$ and task $B$. Here, the accuracy on task $A$ evaluates how well the models retain knowledge of the original task without forgetting, while the accuracy on task $B$ gauges the models’ ability to learn new tasks.
@@
We would like to note that this setup aims to simulate the actual scenario of model incremental development, where we typically have limited access to the training data of task $A$. Therefore, we introduce ‘replay’ as a baseline for comparison, which refers to relaxing the restrictions on accessing the data of task $A$ while learning task $B$. This is done by revisiting some of the task $A$ data to aid the model in remembering and maintaining its performance on the old task. In particular, we aim to compare the following seven methods: naive retraining (i.e., directly training with the data of task B), Replay 5% (i.e., the training data includes 5% of task $A$ data with all data of task $B$), Replay 10%, Replay 100%, DeepArc (i.e., training task $B$ only on the redundant layers of the model), NeuSemSlice, and NeuSemSlice combined with Replay.
@@
Results and Findings. The results, as shown in Table 6, initially reveal that direct retraining causes catastrophic forgetting in every setting, with the model completely forgetting knowledge related to task $A$ in four settings. Replay methods outperform direct retraining, with 5% data replay resulting in an average 30.45% decrease in predictive accuracy for task $A$, 10% data replay leading to a 23.03% average decrease, and 100% data replay causing a 5.15% average decrease.
@@
Our method significantly outperforms the 5% and 10% data replay and even surpasses the 100% data replay in four experimental settings. The average precision loss for task $A$ is only 2.79%, with an average accuracy advantage of 0.69% over the 100% data replay. When combined with 10% data replay, our method consistently achieves the best results in all experimental settings, with an average precision loss for task $A$ of only 0.25% and an average accuracy of 90.76%.
@@
These results thoroughly demonstrate that our method effectively mitigates the issue of catastrophic forgetting in task $A$. Even with a fixed portion of the model’s parameters, our method incurs only around a 1.64% loss in accuracy on task $B$ compared to direct retraining. Notably, in terms of storage efficiency, replay methods require storing some training data from task $A$, typically needing several to tens of megabytes of storage space under our experimental settings. In contrast, our method only needs to retain index information of critical neurons in the model, significantly reducing storage requirements to just a few kilobytes, which is particularly crucial for applications with limited storage resources. Moreover, in terms of efficiency, our method is slightly superior to the 5% and 10% data replay and is twice as efficient as the 100% data replay, with more details provided in (NeuSemSlice, 2024). Overall, considering the combined performance on tasks $A$ and $B$, our method demonstrates significant effectiveness in the model incremental development task.
@@
Regarding DeepArc, we find that training on the identified redundant layers of the model does not effectively alleviate catastrophic forgetting. It maintains only partial capability to complete task $A$ on VGG16 and ResNet101, and DeepArc leads to a substantial reduction in the model’s ability to learn new tasks. A possible explanation is that training only the model’s redundant layers is insufficient for learning new tasks.

### v2

@@ 5. Evaluation
In this section, we evaluate the performance of NeuSemSlice. We first outline the experimental setup and then we introduce our evaluation aiming to answer the following research questions:
@@
  • RQ2: Is NeuSemSlice effective for model restructure?
  • RQ3: Is NeuSemSlice effective for model re-adaptation?
@@ 5.1. Evaluation Setup
To ensure the comprehensiveness and depth of the experiments, we select four widely utilized datasets in the deep neural network domain—MNIST (LeCun et al., 1998), Fashion-MNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009), and ImageNet (Deng et al., 2009)—along with five network models for experimental evaluation: MLP, AlexNet (Krizhevsky et al., 2012), VGG-16 (Simonyan and Zisserman, 2015), ResNet-101 (He et al., 2016), and VGG-19 (Simonyan and Zisserman, 2015). These models and datasets are not only widely used in the deep learning domain but also extensively applied in software engineering research (Li et al., 2024d; Jiang et al., 2024; Wang et al., 2024a; Chen et al., 2020; Huang et al., 2024; Liu et al., 2024; Kang et al., 2024).
MNIST and Fashion-MNIST. For training, we use a Multilayer Perceptron (MLP) model with two hidden layers containing 512 and 256 neurons, respectively.
CIFAR-10. Three deep neural network models—AlexNet, VGG-16, and ResNet-101—are utilized for training on this dataset.
ImageNet. Following the setup in (Xie et al., 2022), we select data from the top 10 categories for experiments, employing the VGG-19 model.
@@
All these experiments are conducted on the Intel Xeon Silver 4214 Processor with 2 Tesla V100 GPUs with 32GB memory. We have implemented our tool based on Pytorch (Paszke et al., 2019). The source code, related information of datasets and models, and complete experimental results are released at (NeuSemSlice, 2025).
@@ 5.2. RQ1: The performance of semantic slicing
Detailed Configurations. In RQ1, we compare the following three slicing methods: global slicing, category-wise slicing, and semantic slicing. To assess neuron contribution and select critical neurons, we utilize five neuron contribution metrics: Avg, Var, Ens, DeepLIFT, and Taylor. As outlined in Section 3.1, global slicing and category-wise slicing choose top 10%, 20%, … up to 100% of neurons in each layer or category for slicing respectively. Additionally, we introduce DeepArc for comparative analysis. Following the settings used in (Ren et al., 2023), we set the threshold at 0.99 to identify redundant layers in the model. Subsequently, we remove these redundant layers and compare the model accuracy when only critical semantic layers are retained.
@@
Comparison with Global Slicing and Category-wise Slicing. Figure 5 compares the slicing precision differences among semantic slicing, global slicing, and category-wise slicing (the complete experimental results can be found in (NeuSemSlice, 2025)). For complex models like AlexNet and VGG16, global slicing and category-wise slicing can only maintain certain model performance at very low compression rates, with accuracy significantly declining as the compression rate increases. Semantic slicing demonstrates significant improvement over the other two techniques across all five neuron contribution metrics. For example, in the AlexNet-CIFAR10 case (Figure 5(c)), semantic slicing uses the neuron contribution metric Taylor to maintain model performance, with only a 0.38% loss in accuracy at a 50% compression rate. Conversely, the other two techniques experience significant accuracy losses of around 70% at the same compression rate. Similarly, on VGG16 using the DeepLIFT metric (Figure 5(d)), semantic slicing at around a 10% compression rate reduces accuracy loss by 41.18% and 77.77% compared to global slicing and category-wise slicing.
@@
Table 4. Statistical Comparison of Semantic Slicing With Global and Category-Wise Slicing Across Various Experimental Settings.
Settings Semantic vs Global Semantic vs Category-wise
$p$-value Cohen’s $d$ $p$-value Cohen’s $d$
MLP-MNIST **∗ ∗ **∗ ∗ * *
MLP-FMNIST ***∗ ∗ ∗ ***∗ ∗ ∗ **∗ ∗ **∗ ∗
AlexNet-CIFAR10 ***∗ ∗ ∗ ***∗ ∗ ∗ ***∗ ∗ ∗ ***∗ ∗ ∗
VGG16-CIFAR10 **∗ ∗ ***∗ ∗ ∗ ***∗ ∗ ∗ ***∗ ∗ ∗
ResNet101-CIFAR10 ***∗ ∗ ∗ ***∗ ∗ ∗ ***∗ ∗ ∗ ***∗ ∗ ∗
VGG19-ImageNet ***∗ ∗ ∗ ***∗ ∗ ∗ ***∗ ∗ ∗ ***∗ ∗ ∗
  • $p$-value: *** ¡ 0.01 (highly significant), ** ¡ 0.05 (significant), * ¿ 0.05 (not significant).
  • Cohen’s $d$: *** (¿ 0.8, large effect), ** (0.5 - 0.8, medium effect), * (0.2 - 0.5, small effect).
To further clarify the extent of the differences among these approaches, we conduct paired $t$-tests and compute effect sizes using Cohen’s $d$ on the AUC values obtained within the same range from the compression rate-accuracy curves (see Table 4). These tests confirm that semantic slicing achieves statistically significant improvements over global slicing in all experimental configurations ($p$-values ¡ 0.05) and, except for the MLP-MNIST setting, also significantly outperforms category-wise slicing. Moreover, the corresponding effect sizes are medium to large in most cases, indicating that the advantages of semantic slicing are not only statistically robust but also practically meaningful.
These results adequately confirm that an inadequate account of the uneven distribution of critical neurons across different categories and network layers leads to poor slicing performance. This highlights the advantage of the proposed technique, which can dynamically identify critical features and components of the model based on semantic similarity.
The Effects of DeepArc. For DeepArc, we identify redundant layers in six settings based on layer similarity, as shown in Figure 5. Firstly, for the MLP, the model only has two hidden layers. When any layer is removed, we follow the model restructuring detailed in DeepArc and introduce a “sewing layer” to align the input and output of the layers. However, this results in no change in the size of the model, indicating that DeepArc cannot effectively identify redundancies in such models. In the other four configurations, DeepArc only identifies redundant layers with a similarity greater than 0.99 in VGG16 and ResNet101.
@@
5.3. RQ2: Model Restructure
Table 5. Results of Model Restructure.
@@ 5.3. RQ2: Model Restructure
Detailed Configuration. To systematically assess NeuSemSlice for the model restructure task, we adopt the same experimental setup detailed in (Ren et al., 2023), and validate by comparing the compression efficiency and effects on the model before and after applying NeuSemSlice. Specifically, after removing the non-critical neurons of the model, we expect the compression technology to directly focus on the redundant parts within the critical neurons, improving the iteration efficiency [^3] It is important to note that in this experiment and the subsequent RQ3 and RQ4, our method is based on the optimal configuration identified in RQ1, while global slicing and category-wise slicing use configurations with similar slice sizes under the same neuron contribution metric.. Based on this, we apply four of the most advanced model compression techniques currently available—LAMP (Lee et al., 2021), Global (Morcos et al., 2019), Uniform+ (Gale et al., 2019), and ERK (Evci et al., 2020)—for optimal compression results. Additionally, we also present the results of global slicing and category-wise slicing for comparison.
@@
Results and Findings. Table 5 presents the experimental results of NeuSemSlice in model restructure. Across all experimental settings, the application of NeuSemSlice in model restructure allows models to achieve higher compression ratios with fewer iterations, indicating that NeuSemSlice can improve both compression efficiency and effectiveness. Apart from the VGG19-ImageNet configuration, NeuSemSlice not only enhances compression efficiency and rates but also slightly improves model accuracy. Specifically, in six experimental settings, the average number of iterations required for model compression decreases from 8.67 to 7.50, enhancing efficiency by 13.46%, and the model compression ratio increases from 37.26% to 40.26%.
For global slicing and category-wise slicing, we observe that in simpler models like MLP-MNIST and MLP-FMNIST, their performance is similar to NeuSemSlice, both effectively improving compression efficiency. However, as the model complexity increases, NeuSemSlice’s ability to accurately identify critical neurons becomes more apparent. In the case of AlexNet-CIFAR10, with the same number of compression iterations and similar compression ratios, NeuSemSlice achieves an average improvement of 1.67% in model accuracy compared to global slicing and a 7.30% improvement compared to category-wise slicing. For more complex models, such as VGG16-CIFAR10, ResNet101-CIFAR10, and VGG19-ImageNet, global slicing and category-wise slicing often result in significant accuracy drops. For instance, global slicing reduces the accuracy of VGG16-CIFAR10 to around 47%, while category-wise slicing drops the accuracy of VGG19-ImageNet to about 58%, whereas NeuSemSlice does not experience such issues. These results demonstrate that the application of NeuSemSlice in model restructure boosts the performance of model compression and enhances the performance of compressed models.
Answer to RQ2: For model restructure, NeuSemSlice not only elevates the efficiency of compression technology—an average increase of 13.46%—but also achieves approximately a 3% enhancement in the compression rate and a slight increase in accuracy.
@@ 5.4. RQ3: Model Re-adaptation
Table 6. Results of Model Re-adaptation.
@@
Method TPR $\downarrow$ RA $\uparrow$ RAA      $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$
@@
Detailed Configuration. To answer RQ3, we conduct an experiment using two prevalently-employed adversarial sample generation techniques, Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015) and Projected Gradient Descent (PGD) (Madry et al., 2018). We generate 1,000 adversarial samples, with FGSM reducing the model’s average accuracy to 18.53%, and PGD to 3.8%. We aim to compare the Training Parameter Ratio (TPR) and the Retrained Accuracy (RA) of existing approaches and our proposed method. Specifically, the experiment settings include fully training (i.e., comparison baseline), DeepArc, global slicing, category-wise slicing, NeuSemSlice, and a combination of NeuSemSlice with DeepArc [^4] The term “a combination of NeuSemSlice with DeepArc” refers to retraining the critical neurons identified by our method in the critical modules recognized by DeepArc. under identical training epochs. Additionally, this RQ further assesses the resilience of each method against similar adversarial attacks. We generate another 1,000 adversarial samples after the model re-adaptation, and obtain the model accuracy against the samples (i.e., shown as Re-Attack Accuracy, RAA).
Results and Findings. Table 6 presents our experimental results in the six settings. Compared to the traditional method of fully training and DeepArc, our method demonstrates a significant reduction in the required model parameters for training, averaging only 61.49%. This represents an average decrease of 38.51% in training parameters compared to the fully training strategy and an average reduction of 16.66% relative to DeepArc. Moreover, our method surpasses fully training by an average of 0.67% and DeepArc by an average of 0.50% in Retrained Accuracy, demonstrating an advantage in both training cost-efficiency and effectiveness. Compared to global slicing and category-wise slicing, with similar training parameters, our method shows an average improvement of 2.91% and 2.89% in Retrained Accuracy, respectively. This highlights that the inaccurate identification of critical neurons leads to reduced Re-adaptation effectiveness. Although fewer training parameters are used than the fully training method, the training accuracy also decreases accordingly. Among the six methods, our method combined with DeepArc achieves a 0.73% increase in Retrained Accuracy compared to traditional training methods, with the lowest training cost (averaging only 47.75% of the parameters trained).
Our method combined with DeepArc also shows the best performance in terms of re-attack performance, achieving up to a 17.8% improvement in RAA. The results further demonstrate the superior performance of our method in model re-adaptation scenarios. However, we observe that all of the methods experience notable performance loss against re-attack, alerting us to design better methods against adversarial attacks in the future, to be discussed in Section 5.7. Finally, NeuSemSlice is comparable to DeepArc in efficiency, requiring only 93.53% of the time needed for full training on average. When NeuSemSlice is combined with DeepArc, it requires only 89.80% of the full training time. The detailed data is available at (NeuSemSlice, 2025).
@@ 5.5. RQ4: Model Incremental Development
Table 7. Results of Model Incremental Development(Scenario I)
@@
Table 8. Results of Model Incremental Development(Scenario II)
Retrain Replay 5% Replay 10% Global Category-wise NeuSemSlice
MLP-MNIST TaskA 0.00% 91.93% 94.85% 94.22% 93.55% 98.95%
TaskB 0.14% 85.89% 90.71% 97.54% 98.79% 97.43%
TaskC 98.11% 97.96% 97.88% 97.88% 97.98% 97.93%
Avg 32.75% 91.93% 94.48% 96.55% 96.77% 98.11%
MLP-FMNIST TaskA 0.00% 58.93% 66.40% 84.90% 90.13% 93.57%
TaskB 0.00% 55.47% 68.63% 91.03% 96.10% 96.73%
TaskC 96.80% 95.58% 94.63% 96.15% 96.25% 96.15%
Avg 32.27% 69.99% 76.55% 90.69% 94.16% 95.48%
AlexNet-CIFAR10 TaskA 0.00% 37.57% 57.17% 75.57% 87.83% 93.37%
TaskB 0.00% 29.30% 50.83% 59.07% 70.17% 76.03%
TaskC 94.63% 94.60% 93.63% 92.70% 94.10% 92.93%
Avg 31.54% 53.82% 67.21% 75.78% 84.03% 87.44%
VGG16-CIFAR10 TaskA 0.00% 38.63% 54.77% 86.90% 76.57% 89.17%
TaskB 0.00% 33.10% 45.67% 54.33% 60.80% 65.93%
TaskC 94.53% 92.30% 91.30% 91.40% 89.65% 91.55%
Avg 31.51% 54.68% 63.91% 77.54% 75.67% 82.22%
ResNet101-CIFAR10 TaskA 10.03% 17.90% 38.37% 91.90% 92.20% 92.30%
TaskB 6.37% 15.70% 27.63% 64.60% 57.07% 62.63%
TaskC 89.73% 90.70% 87.13% 86.80% 89.73% 89.35%
Avg 35.38% 41.43% 51.04% 81.10% 79.66% 81.43%
VGG19-
ImageNet
TaskA 0.00% 31.28% 32.44% 93.33% 92.31% 98.21%
TaskB 0.00% 28.08% 29.74% 80.13% 83.72% 85.13%
TaskC 92.31% 92.02% 91.44% 93.08% 93.46% 92.60%
Avg 30.77% 50.46% 51.21% 88.85% 89.83% 91.98%
@@
Detailed Configuration. To answer RQ4, we set up two experimental scenarios. In the first scenario, we split the datasets MNIST, FMNIST, CIFAR10, and ImageNet into two subtasks, each containing five categories, labeled as task $A$ (previous task) and task $B$ (new task). For example, in the CIFAR10 dataset, which consists of ten distinct image categories, task A includes the first five categories, while task B includes the remaining five categories. In the second scenario, we further divide the datasets into three subtasks. For instance, task $A$ and task $B$ include three categories each, while task $C$ includes four categories.
@@
To simulate the process of incremental model development, for the first scenario, we start by training models exclusively on task $A$ in six different experimental setups following the division. Then, we proceed to train these models on task $B$ while aiming to preserve their performance on task $A$. The evaluation focuses on the accuracy for both task $A$ and task $B$. Here, the accuracy on task $A$ evaluates how well the models retain knowledge of the original task without forgetting, while the accuracy on task $B$ gauges the models’ ability to learn new tasks. The second scenario follows a similar process (with an additional round of new task learning). Here, the model first learns task A, followed by task B, and finally task C.
@@
We would like to note that this setup aims to simulate the actual scenario of model incremental development, where we typically have limited access to the training data of previous tasks. Therefore, we introduce ‘replay’ as a baseline for comparison, which refers to relaxing the restrictions on accessing the data of previous tasks while learning a new task. This is done by revisiting some of the task $A$ data to aid the model in remembering and maintaining its performance on the old task.
In particular, we aim to compare the following methods: naive retraining (i.e., directly training with new task data only), Replay 5% (i.e., training data includes 5% of previous task data along with all new task data), Replay 10%, Replay 100%, DeepArc (i.e., training the new task only on the redundant layers of the model), global slicing, category-wise slicing, NeuSemSlice, and NeuSemSlice combined with Replay.
@@
Results and Findings. The results of Experiment Scenario 1, as shown in Table 7, indicate that direct retraining causes catastrophic forgetting in all settings, with the model completely forgetting knowledge related to task $A$ in six settings. Replay methods outperform direct retraining, with 5% data replay resulting in an average 30.45% decrease in predictive accuracy for task $A$, 10% data replay leading to a 23.03% average decrease, and 100% data replay causing a 5.15% average decrease. Additionally, global slicing and category-wise slicing also achieve promising results, with global slicing reducing task $A$’s accuracy by 11.75% and category-wise slicing by 10.58%, both outperforming 5% and 10% data replay.
@@
Our method outperforms 5% and 10% data replay, as well as global slicing and category-wise slicing, and achieves comparable performance to 100% data replay in most scenarios, with an average precision loss of only 2.79% on task $A$. When combined with 10% data replay, our method consistently achieves the best results in all experimental settings, with an average precision loss for task $A$ of only 0.25% and an average accuracy of 90.76%.
@@
These results thoroughly demonstrate that our method effectively mitigates the issue of catastrophic forgetting in task $A$. Even with a fixed portion of the model’s parameters, our method incurs only around a 1.64% loss in accuracy on task $B$ compared to direct retraining. Notably, in terms of storage efficiency, replay methods require storing some training data from task $A$, typically needing several to tens of megabytes of storage space under our experimental settings. In contrast, our method only needs to retain index information of critical neurons in the model, significantly reducing storage requirements to just a few kilobytes, which is particularly crucial for applications with limited storage resources. Moreover, in terms of efficiency, our method is slightly superior to the 5% and 10% data replay and is twice as efficient as the 100% data replay, with more details provided in (NeuSemSlice, 2025). Overall, considering the combined performance on tasks $A$ and $B$, our method demonstrates significant effectiveness in the model incremental development task.
@@
Regarding DeepArc, we find that training on the identified redundant layers of the model does not effectively alleviate catastrophic forgetting. It maintains only partial capability to complete task $A$ on VGG16 and ResNet101, and DeepArc leads to a substantial reduction in the model’s ability to learn new tasks.
The results for the second scenario appear in Table 8. We further compare NeuSemSlice with naive retraining, replay-based methods, global slicing, and category-wise slicing  [^5] Due to the significant decline in DeepArc’s ability to learn new tasks observed in Experiment Scenario I, it is unable to proceed with training task C. Therefore, we do not include it in the comparison here.. In six experimental settings, direct retraining once again forgets knowledge related to task $A$ and task $B$. Replay-based methods also decline compared to the first scenario. For example, replay that uses 5% of the data yields average accuracies of 91.93%, 69.99%, 53.82%, 54.68%, 41.43%, and 50.46% across the six settings, while replay that uses 10% of the data produces similar results and performs well only on MLP-MNIST. In contrast, NeuSemSlice achieves the highest performance, with an overall average accuracy of 89.44% in the six settings, which is 4.36% higher than global slicing and 2.75% higher than category-wise slicing. However, the introduction of more complex tasks leads to imbalanced performance across tasks, with forgetting in task B being the most severe in most settings. A possible explanation is that task A, as the base model, is deeply embedded in the model’s parameters, while task C receives focused training as the newest target, leading to stronger performance on these two tasks.
To further validate the effectiveness of NeuSemSlice, we perform paired $t$-tests across all experimental settings to compare its performance with baseline methods, including replay-based methods, global slicing, and category-wise slicing. The results show that NeuSemSlice’s improvements are statistically significant ($p$-values ¡ 0.05, Cohen’s $d$ 0.8) in all comparisons except for replay 10% ($p$-values = 0.06, Cohen’s $d$ = 0.98) and replay 100% ($p$-values = 0.33, Cohen’s $d$ = 0.44) in Experiment Scenario 1. These statistical findings confirm the robustness and generalizability of NeuSemSlice in incremental development scenarios, providing strong evidence of its practical utility in mitigating catastrophic forgetting while maintaining computational and storage efficiency.